OpenStudy (anonymous):

Solve. 1 = –(–w + 6) A. 3 B. –3 C. 7 D. –7

OpenStudy (anonymous):

Distribute the -1 over -w and +6 what happens? Then isolate w.

OpenStudy (anonymous):



OpenStudy (anonymous):

I am very sorry I am on a hurry so could you just give me answer i will medal and fan

OpenStudy (anonymous):

No sorry

OpenStudy (anonymous):

1=-(-w+6) 1=-1(-w+6) 1=w-6 7=w

OpenStudy (anonymous):

C

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

You should really learn the subject, rather than have someone give you the answer, as it will not help you in the future, but as always do what you like, good luck!

OpenStudy (anonymous):

i am sorry im just on a hurry

OpenStudy (anonymous):

Am I good

OpenStudy (anonymous):

No worries :) but if you need help you can always tag me, if you're interested in steps I can show you then.